OpenStudy (anonymous):

classify the solutions of 3/x+5/15=2/x+5 as extraneous or non extranous

OpenStudy (mertsj):

\[\frac{3}{\frac{x+5}{15}}=\frac{2}{x+5}\]

OpenStudy (mertsj):

Is that the problem?

OpenStudy (sweetburger):

pretty sure its non extranous because no x's make the denominator equal 0

OpenStudy (anonymous):

no there all side buy side

OpenStudy (nikato):

?